### v1

# Two tales for a geometric Jensen–Shannon divergence

Abstract
The geometric Jensen–Shannon divergence (G-JSD) gained popularity in machine learning and information sciences thanks to its closed-form expression between Gaussian distributions. In this work, we introduce an alternative definition of the geometric Jensen–Shannon divergence tailored to positive densities which does not normalize geometric mixtures. This novel divergence is termed the extended G-JSD as it extends to more general positive measures. We give explicitly the gap between the extended G-JSD and G-JSD when considering probability densities, and report both lower and upper bounds in terms of other statistical divergences. We derive corresponding closed-form expressions when considering the case of multivariate Gaussian distributions often met in applications. Finally, we show that these two types of geometric JSDs, the G-JSD and the extended G-JSD, can be interpreted as regularizations of the ordinary JSD by additive terms.
Keywords: Jensen–Shannon divergence; power means; Chernoff information; Bhattacharyya distance; geometric mixtures; exponential families.
1 Introduction
Let $(\mathcal{X},\mathcal{E},\mu)$ be a measure space on the sample space $\mathcal{X}$, $\sigma$-algebra of events $\mathcal{E}$, with $\mu$ a prescribed positive measure on the measurable space $(\mathcal{X},\mathcal{E})$ (e.g., counting measure or Lebesgue measure). Let $M_{+}(\mathcal{X})=\{Q\}$ be the set of positive distributions $Q$ and $M_{+}^{1}(\mathcal{X})=\{P\}$ be the subset of probability measures $P$. We denote by $M_{\mu}=\{\frac{\mathrm{d}Q}{{\mathrm{d}\mu}}{\ :\ }Q\in M_{+}(\mathcal{X})\}$ and $M^{1}_{\mu}=\{\frac{\mathrm{d}P}{{\mathrm{d}\mu}}{\ :\ }P\in M_{+}^{1}(\mathcal{X})\}$ the corresponding sets of Radon-Nikodym positive and probability densities, respectively.
Consider two probability measures $P_{1}$ and $P_{2}$ of $M_{+}^{1}(\mathcal{X})$ with Radon-Nikodym densities with respect to $\mu$ $p_{1}{:=}\frac{\mathrm{d}P_{1}}{{\mathrm{d}\mu}}\in M^{1}_{\mu}$ and $p_{2}{:=}\frac{\mathrm{d}P_{2}}{{\mathrm{d}\mu}}\in M_{\mu}^{1}$, respectively. The deviation of $P_{1}$ to $P_{2}$ (also called distortion, dissimilarity, or deviance) is commonly measured in information theory [6] by the Kullback–Leibler divergence (KLD):
$D_{\mathrm{KL}}(p_{1},p_{2}){:=}\int p_{1}\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}=E_{p_{1}}\left[\log\frac{p_{1}}{p_{2}}\right].$ (1)
Informally, the KLD quantifies the information lost when $p_{2}$ is used to approximate $p_{1}$ by measuring on average the surprise [^1] Shannon entropy $H(p)=\int p\log\frac{1}{p}\,{\mathrm{d}\mu}$ is the expected surprise $H(p)=E_{p}[-\log p]$ where $\log p(x)$ measures the surprise of the outcome $x$. when outcomes sampled from $p_{1}$ are assumed to emanate from $p_{2}$. Gibbs’ inequality assert that $D_{\mathrm{KL}}(P_{1},P_{2})\geq 0$ if equality if and only if $P_{1}=P_{2}$ $\mu$-almost everywhere. Since $D_{\mathrm{KL}}(p_{1},p_{2})\not=D_{\mathrm{KL}}(p_{2},p_{1})$, various symmetrization schemes of the KLD have been proposed in the literature [6] (e.g., Jeffreys divergence [6, 14], resistor average divergence [15] (harmonic symmetrization), Chernoff information [6], etc.)
An important symmetrization technique of the KLD is the Jensen–Shannon divergence [19, 10] (JSD):
$D_{\mathrm{JS}}(p_{1},p_{2}):=\frac{1}{2}\,\left(D_{\mathrm{KL}}(p_{1},a)+D_{\mathrm{KL}}(p_{2},a)\right),$ (2)
where $a=\frac{1}{2}p_{1}+\frac{1}{2}p_{2}$ denotes the statistical mixture of $p_{1}$ and $p_{2}$. The JSD is guaranteed to be upper bounded by $\log 2$ even when the support of $p_{1}$ and $p_{2}$ differ, making it attractive in applications. Furthermore, its square root $\sqrt{D_{\mathrm{JS}}}$ yields a metric distance [9, 31]. The JSD can be extended to a set of densities to measure the diversity of the set as an information radius [35]. In information theory, the JSD can also be interpreted as an information gain [9] since it can be equivalently written as
$D_{\mathrm{JS}}(p_{1},p_{2})=H\left(\frac{1}{2}p_{1}+\frac{1}{2}p_{2}\right)-\frac{H(p_{1})+H(p_{2})}{2},$
where $H(p)=-\int p\log p,{\mathrm{d}\mu}$ is Shannon entropy. [^2] Shannon entropy for discrete measures and differential entropy for continuous measures. The JSD was also defined in the setting of quantum information [5] where it was also proven that its square root yields a metric distance [38].
In [25], the following generalization of the KLD Jensen–Shannon symmetrization scheme was studied for arbitrary statistical dissimilarity $D(\cdot,\cdot)$:
Definition 1 ($(\alpha,\beta)$ M-JS dissimilarity [25])
The Jensen–Shannon skew symmetrization of a statistical dissimilarity $D(\cdot,\cdot)$ with respect to an arbitrary weighted bivariate mean $M_{\alpha}(\cdot,\cdot)$ is given by:
$\mathrm{JS}_{D}^{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{\alpha}^{M}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{\alpha}^{M}\right),\quad(\alpha,\beta)\in(0,1)^{2},$ (3)
where ${(p_{1}p_{2})}_{\alpha}^{M}$ is the statistical normalized $M$-mixture of $p_{1}$ and $p_{2}$:
${(p_{1}p_{2})}_{\alpha}^{M}(x){:=}\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)}.$ (4)
(A more general definition is given in [25] by using another arbitrary weighted mean $N_{\beta}$ to average the two KLDs in Eq. 3).
To avoid burden of notations, we write for short $\mathrm{JS}_{D}^{M}(p_{1},p_{2})$ when $\alpha=\beta=\frac{1}{2}$, and further let $D_{\mathrm{JS}}^{M}=\mathrm{JS}_{\mathrm{KL}}^{M}=\mathrm{JS}_{{\mathrm{KL}}}^{M_{\frac{1}{2}},\frac{1}{2}}$.
The ordinary Jensen–Shannon divergence is recovered when $D=D_{\mathrm{KL}}$, $M_{\alpha}(a,b)=\alpha a+(1-\alpha)b$ (weighted arithmetic mean $A_{\alpha}$), and $\beta=\alpha=\frac{1}{2}$. Lower and upper bounds for the skewed $\alpha$-Jensen–Shannon divergence $\mathrm{JS}_{\mathrm{KL}}^{A_{1-\alpha},\frac{1}{2}}$ were reported in [40]. The abstract mixture normalizer of ${(p_{1}p_{2})}_{\alpha}^{M}$ shall be denoted by $Z_{\alpha}^{M}(p_{1},p_{2}){:=}\int M_{\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}(x)$ so that the normalized $M$-mixture is ${(p_{1}p_{2})}_{\alpha}^{M}(x)=\frac{M_{\alpha}(p_{1}(x),p_{2}(x))}{Z_{\alpha}^{M}(p_{1},p_{2})}$.
The generic Jensen–Shannon symmetrization of Definition 1 allows to re-interpret some well-known statistical dissimilarities: For example, the Chernoff information [6, 26] defined by
$D_{C}(p_{1},p_{2}){:=}\max_{\alpha\in(0,1)}D_{B,\alpha}(p_{1},p_{2}),$ (5)
with $D_{B,\alpha}(p_{1},p_{2}){:=}-\log\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$ the $\alpha$-skewed Bhattacharrya divergence is a $(\frac{1}{2},\alpha^{*})$ G-$D_{\mathrm{KL}}^{*}$ Jensen–Shannon divergence for the geometric mean $G_{\alpha}(a,b)=a^{\alpha}b^{1-\alpha}$ since for the provably unique optimal skewing factor $\alpha^{*}$ in Eq. 5, we have [26]:
$D_{C}(p_{1},p_{2})=D_{\mathrm{KL}}^{*}(p_{1},(p_{1}p_{2})_{\alpha^{*}}^{G})=D_{\mathrm{KL}}^{*}(p_{2},(p_{1}p_{2})_{\alpha^{*}}^{G}),$
where $D_{\mathrm{KL}}$$^{*}$ is the reverse KLD: $D_{\mathrm{KL}}^{*}(p_{1},p_{2}){:=}D_{\mathrm{KL}}(p_{2},p_{1})$. Hence, it follows that $D_{C}(p_{1},p_{2})=\mathrm{JS}_{{D_{\mathrm{KL}}^{*}}}^{G_{\alpha^{*}},\frac{1}{2}}(p_{1},p_{2})$.
Although arithmetic mixtures are most often used in Statistics, the geometric mixtures are also encountered, like for example in Bayesian statistics [2], or in Markov chain Monte Carlo annealing [12], just to give two examples. In information geometry, statistical power mixtures based on the homogeneous power means are used to perform stochastic integration of statistical models [1].
Proposition 1 (G-JSD as Bhattacharyya distance)
We have
$\mathrm{JS}_{{\mathrm{KL}}*}^{G}(p_{1},p_{2})=-\log\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}=:D_{B}(p_{1},p_{2}),$
where $D_{B}$ denotes the Bhattacharyya distance [4], and
$\mathrm{JS}_{{\mathrm{KL}}*}^{{G_{\alpha}},\alpha}(p_{1},p_{2})=-\log\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}=:D_{B,\alpha}(p_{1},p_{2}),$
whee $D_{B,\alpha}$ denotes the $\alpha$-skewed Bhattacharyya distance.
Proof:
Let $m=(p_{1}p_{2})^{G}=\frac{\sqrt{p_{1}p_{2}}}{Z(p_{1},p_{2})}$ be the weighted geometric mixture with normalizer $Z(p_{1},p_{2})=\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}$. By definition of the JS-symmetrization of the reverse KLD, we have
$\mathrm{JS}_{{\mathrm{KL}}*}^{G}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left(D_{{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})^{G})+D_{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})^{G})\right),$
$=$ $\frac{1}{2}\left(D_{{\mathrm{KL}}}((p_{1}p_{2})^{G},p_{1},)+D_{{\mathrm{KL}}}((p_{1}p_{2})^{G},p_{2},)\right),$
$=$ $\frac{1}{2}\left(m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{1}Z(p_{1},p_{2})}+m\,\log\frac{\sqrt{p_{1}p_{2}}}{p_{2}Z(p_{1},p_{2})}\right),$
$=$ $\frac{1}{2}\left(\int\frac{1}{2}m\log\frac{p_{2}}{p_{1}}\,\frac{p_{1}}{p_{2}}{\mathrm{d}\mu}-2\log Z(p_{1},p_{2})\int m\,{\mathrm{d}\mu}\right),$
$=$ $-\log Z(p_{1},p_{2}).$
The proof carries on similarly for the $\alpha$-skewed JS-symmetrization of the reverse KLD: We now let $m_{\alpha}=(p_{1}p_{2})^{G}_{\alpha}=\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}(p_{1},p_{2})}$ be the $\alpha$-weighted geometric mixture with normalizer $Z_{\alpha}(p_{1},p_{2})=\int p_{1}^{\alpha}p_{2}^{1-\alpha}\,{\mathrm{d}\mu}$, written as $Z_{\alpha}$ for short below:
$\mathrm{JS}_{{\mathrm{KL}}*}^{G_{\alpha},\alpha}(p_{1},p_{2})$ ${:=}$ $\alpha\,D_{{\mathrm{KL}}^{*}}(p_{1},(p_{1}p_{2})^{G}_{\alpha})+(1-\alpha)\,D_{{\mathrm{KL}}^{*}}(p_{2},(p_{1}p_{2})^{G}_{\alpha}),$
$=$ $\alpha\,D_{{\mathrm{KL}}}(m_{\alpha},p_{1})+(1-\alpha)\,D_{{\mathrm{KL}}}(m_{\alpha},p_{2}),$
$=$ $\int\left(\alpha m_{\alpha}\log\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}p_{1}}+(1-\alpha)m_{\alpha}\log\frac{p_{1}^{\alpha}p_{2}^{1-\alpha}}{Z_{\alpha}p_{2}}\right){\mathrm{d}\mu},$
$=$ $-(\alpha+1-\alpha)\log Z_{\alpha}\int m_{\alpha}\,{\mathrm{d}\mu}+\int m_{\alpha}\log\left(\frac{p_{2}}{p_{1}}\right)^{\alpha(1-\alpha)}\,\left(\frac{p_{1}}{p_{2}}\right)^{\alpha(1-\alpha)}{\mathrm{d}\mu},$
$=$ $-\log Z_{\alpha}(p_{1},p_{2}):=D_{B,\alpha}(p_{1},p_{2}).$
$\square$
Besides information theory, the JSD also plays an important role in machine learning [21, 11, 36]. However, one drawback that refrains its use in practice is that the JSD between two Gaussian distributions is not known in closed-form since no analytic formula is known for the differential entropy of a two-component Gaussian mixture [22], and thus the JSD needs to be numerically approximated in practice by various methods.
Remark 1
In general, the log-mixture term integrand makes it difficult to compute in closed-form the differential entropy of mixture of densities although some exceptions may exist (e.g., the JSD between Cauchy distributions is available in closed-form [29]).
The geometric G-JSD was defined in [25] as follows:
Definition 2 (G-JSD)
The geometric Jensen–Shannon divergence (G-JSD) between two probability densities $p_{1}$ and $p_{2}$ is defined by
$D_{\mathrm{JS}}^{G}(p_{1},p_{2}){:=}\frac{1}{2}\,\left(D_{\mathrm{KL}}(p_{1}:(p_{1}p_{2})^{G})+D_{\mathrm{KL}}(p_{2}:(p_{1}p_{2})^{G})\right),$
where $(p_{1}p_{2})^{G}(x)=\frac{\sqrt{p_{1}(x)p_{2}(x)}}{\int\sqrt{p_{1}(x)p_{2}(x)}\,{\mathrm{d}\mu}}$ is the (normalized) geometric mixture of $p_{1}$ and $p_{2}$.
To contrast with the approximation limitation of the JSD between Gaussians, one advantage of the geometric Jensen–Shannon divergence (G-JSD) is that it admits a closed-form expression between Gaussian distributions [25]. However, the G-JSD is not anymore bounded and thus requires matching support of the densities. The G-JSD formula between Gaussian distributions has been used in several scenarii. See [8, 7, 20, 18, 23, 33, 39, 34, 37, 13]) for a few use cases.
The paper is organized as follows: We first give an alternative definition of the M-JSD in §2 (Definition 3) which extends to positive measures and do not require normalization of geometric mixtures. We call this new divergence the extended M-JSD and we compare the two types of geometric JSDs when dealing with probability measures. In §4, we show that these normalized/extended $M$-JSD divergences can be interpreted as regularizations of the Jensen–Shannon divergence, and exhibit several bounds. We report related closed-form formula for the G-JSD and extended G-JSD between two Gaussian distributions in Section 3. Finally, we summarize the main results in the concluding section §5.
2 A novel definition G-JSD extended to positive measures
We may consider the following two modifications of the G-JSD:
  • First, we replace the KLD by the extended KLD (see [30] for more general extended $f$-divergences) between positive densities $q_{1}\in M_{\mu}^{+}$ and $q_{2}\in M_{\mu}^{+}$ instead of normalized densities:
    $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}\int\left(q_{1}\log\frac{q_{1}}{q_{2}}+q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$ (6)
    and,
  • Second, we consider unnormalized $M$-mixture densities ${(q_{1}q_{2})}_{\alpha}^{\tilde{M}}(x){:=}{M_{\alpha}(q_{1}(x),q_{2}(x))}$ where we use the $\tilde{M}$ notation to indicate that the $M$-mixture is not normalized, instead of normalized densities ${(q_{1}q_{2})}_{\alpha}^{M}(x)$.
The extended KLD can be interpreted as a pointwise integral of a scalar Bregman divergence obtained for the negative Shannon entropy generator [16]. This proves that $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})\geq 0$ with equality if and only if $q_{1}=q_{2}$ $\mu$-almost everywhere. Notice that $D_{\mathrm{KL}}(q_{1},q_{2})$ may be negative when $q_{1}$ and/or $q_{2}$ are not normalized to probability densities but we always have $D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})\geq 0$.
Remark 2
As a side remark, it is preferable in practice to estimate the KLD between $p_{1}$ and $p_{2}$ by Monte Carlo methods using Eq. 6 instead of Eq. 1 in order to guarantee the non-negativeness of the KLD (Gibbs’ inequality). Indeed, the sampling of $s$ samples $x_{1},\ldots,x_{s}$, defines two unnormalized distributions $q_{1}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{1}(x)\delta_{x_{i}}(x)$ and $q_{2}(x)=\frac{1}{s}\sum_{i=1}^{s}p_{2}(x)\delta_{x_{i}}(x)$ where $\delta_{x_{i}}(x)=\left\{\begin{array}[]{ll}1,&\mbox{if $x=x_{i}$}\cr 0,&\mbox{otherwise}\end{array}\right$..
Remark 3
For an arbitrary distortion measure $D^{+}(q_{1},q_{2})$ between positive measures $q_{1}$ and $q_{2}$, we can build a corresponding projective divergence $\tilde{D}(q_{1},q_{2})$ as follows:
$\tilde{D}(q_{1},q_{2}){:=}D^{+}\left(\frac{q_{1}}{Z(q_{1})},\frac{q_{1}}{Z(q_{2})}\right),$
where $Z(q){:=}\int q\,{\mathrm{d}\mu}$ is the normalization factor of the positive density $q$. The divergence $\tilde{D}$ is said projective because we have for all $\lambda_{1}>0,\lambda_{2}>0$, the property that $\tilde{D}(\lambda_{1}q_{1},\lambda_{2}q_{2})=\tilde{D}(q_{1},q_{2})=D^{+}(p_{1},p_{2})$ where $p_{i}=\frac{q_{i}}{Z(q_{i})}$ are the normalized densities. The projective Kullback–Leibler divergence $\tilde{D}_{\mathrm{KL}}$ is thus another projective extension of the KLD to non-normalized densities which coincide with the KLD for probability densities. But the projective KLD is different from the extended KLD of Eq. 6 and furthermore we have $\tilde{D}_{\mathrm{KL}}(q_{1},q_{2})=0$ if and only if $q_{1}=\lambda q_{2}$ $\mu$-almost everywhere for some $\lambda>0$.
Let us now define the Jensen–Shannon symmetrization of an extended statistical divergence $D^{+}$ with respect to an arbitrary weighted mean $M_{\alpha}$ as follows:
Definition 3 (Extended M-JSD)
A Jensen–Shannon skew symmetrization of a statistical divergence $D^{+}(\cdot,\cdot)$ between two positive measures $q_{1}$ and $q_{2}$ with respect to a weighted mean $M_{\alpha}$ is defined by
[…]
When $\beta=\frac{1}{2}$, we simplify the notation $\mathrm{JS}_{D^{+}}^{\tilde{M}_{\alpha}}$ for short, and furthermore when $\alpha=\frac{1}{2}$, we write $\mathrm{JS}_{D^{+}}^{\tilde{M}}$ to avoid burden notations. Let $D_{\mathrm{JS}^{+}}^{M}=\mathrm{JS}_{\mathrm{JS}^{+}}^{\tilde{M}_{\frac{1}{2}},\frac{1}{2}}$.
By abuse of notations, we have
$D_{{\mathrm{KL}}^{+}}(q_{1},q_{2}){:=}D_{\mathrm{KL}}(q_{1},q_{2})+\int\left(q_{2}-q_{1}\right)\,{\mathrm{d}\mu},$
although $q_{1}$ and $q_{2}$ may not need to be normalized in the $D_{\mathrm{KL}}$ term (which can then be potentially negative). Letting $Z(q_{i}){:=}\int q_{i}\,{\mathrm{d}\mu}$ be the total mass of $q_{i}$, we have
$D_{{\mathrm{KL}}^{+}}(q_{1},q_{2})=D_{\mathrm{KL}}(q_{1},q_{2})+Z(q_{2})-Z(q_{1}).$ (8)
Let ${\tilde{m}}_{\alpha}=M_{\alpha}(q_{1},q_{2})$ be the unnormalized $M$-mixture of positive densities $q_{1}$ and $q_{2}$, and set $Z_{\alpha}^{M}=\int{\tilde{m}}_{\alpha}\,{\mathrm{d}\mu}$ be the normalization term so that we have $m_{\alpha}=\frac{{\tilde{m}}_{\alpha}}{Z_{\alpha}}$ and ${\tilde{m}}_{\alpha}=Z_{\alpha}m_{\alpha}$. When clear from context, we write $Z_{\alpha}$ instead of $Z_{\alpha}^{M}$.
We get after elementary calculus:
$\mathrm{JS}_{D_{{\mathrm{KL}}^{+}}}^{\tilde{M}_{\alpha},\beta}(q_{1},q_{2})=\mathrm{JS}_{D_{{\mathrm{KL}}}}^{{M}_{\alpha},\beta}(q_{1},q_{2})-(\beta Z(q_{1})+(1-\beta)Z(q_{2}))\log Z_{\alpha}+Z_{\alpha}-(\beta Z(q_{1})+(1-\beta)Z(q_{2})).$ (9)
Therefore the difference gap $\Delta_{M_{\alpha},\beta}(p_{1},p_{2})$ (written for short as $\Delta(p_{1},p_{2})$) between the normalized JSD and the unnormalized JSD between two normalized densities $p_{1}$ and $p_{2}$ (i.e., with $Z_{1}=Z(p_{1})=1$ and $Z_{2}=Z(p_{2})=1$) is
$\Delta(p_{1},p_{2}){:=}D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})-D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})=Z_{\alpha}-\log(Z_{\alpha})-1.$ (10)
Proposition 2 (Extended/normalize G-JSD Gap)
We have $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1$.
Thus $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\geq 0$ and $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})\leq D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})$ when $\Delta(p_{1},p_{2})\leq 0$.
When we consider the weighted arithmetic mean $A_{\alpha}$, we always have $Z_{\alpha}=1$ for $\alpha\in(0,1)$, and thus the two definitions (Definition 1 and Definition 3) of the $A$-JSD coincide (i.e., $Z_{\alpha}^{A}-\log(Z_{\alpha}^{A})-1=0$):
$D_{\mathrm{JS}}^{A}(p_{1},p_{2})=D_{\mathrm{JS}}^{\tilde{A}}(p_{1},p_{2}).$
However, when the weighted mean $M_{\alpha}$ differs from the weighted arithmetic mean (i.e., $M_{\alpha}\not=A_{\alpha}$), the two definitions of the M-JSD $D_{\mathrm{JS}}^{M}$ and extended M-JSD $D_{\mathrm{JS}}^{\tilde{M}}$ differ by the gap expressed in Eq. 10.
In general, we have $Z_{\alpha}>\log(Z_{\alpha})$ for $Z_{\alpha}>0$, and $Z_{\alpha}-\log(Z_{\alpha})-1<0$ for $Z_{\alpha}<1$ and $Z_{\alpha}-\log(Z_{\alpha})-1\geq 0$ for $Z_{\alpha}\geq 1$.
2.1 Power JSDs and (extended) min-JSD and max-JSD
Let $P_{\gamma,\alpha}(a,b){:=}(\alpha a^{\gamma}+(1-\alpha)b^{\gamma})^{\frac{1}{\gamma}}$ be the $\gamma$-power mean for $\gamma\not=0$ (with $A_{\alpha}=P_{1,\alpha}$). Further define $P_{0,\alpha}(a,b)=G_{\alpha}(a,b)$ so that $P_{\gamma,\alpha}$ defines the weighted power means for $\gamma\in\mathbb{R}$ and $\alpha\in(0,1)$ in the reminder. Since $P_{\gamma,\alpha}(a,b)\leq P_{\gamma^{\prime},\alpha}(a,b)$ when $\gamma^{\prime}\geq\gamma$ for any $a,b>0$, we have that
$Z_{\alpha}^{P_{\gamma}}(p_{1},p_{2})=\int P_{\gamma,\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}\leq Z_{\alpha}^{P_{\gamma^{\prime}}}(p_{1},p_{2})=\int P_{\gamma^{\prime},\alpha}(p_{1}(x),p_{2}(x))\,{\mathrm{d}\mu}.$ (11)
Let us express the G-JSD divergence using other familiar divergences.
Proposition 3
We have the following expression of the geometric Jensen–Shannon divergence:
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}\,D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}),$
where $D_{J}(p_{1},p_{2})=\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [14, 28] and $D_{B}(p_{1},p_{2})=-\log\int\sqrt{p_{1}p_{2}}\,{\mathrm{d}\mu}=-\log Z^{G}(p_{1},p_{2})$ is the Bhattacharrya distance. [^3] The Bhattacharrya distance is not a metric distance as it violates the triangle inequality of metrics.
Proof:
We have:
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})$ $:=$ $\frac{1}{2}\left(D_{\mathrm{KL}}(p_{1},(p_{1}p_{2})^{G})+D_{\mathrm{KL}}(p_{2},(p_{1}p_{2})^{G})\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)\log\frac{p_{1}(x)\,Z^{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)p_{2}(x)}}+p_{2}(x)\log\frac{p_{2}(x)\,Z^{G}(p_{1},p_{2})}{\sqrt{p_{1}(x)p_{2}(x)}}\right){\mathrm{d}\mu}(x)\right),$
$=$ $\frac{1}{2}\left(\int\left(p_{1}(x)+p_{2}(x)\right)\log Z^{G}(p_{1},p_{2})\,{\mathrm{d}\mu}(x)+\frac{1}{2}D_{\mathrm{KL}}(p_{1},p_{2})+\frac{1}{2}D_{\mathrm{KL}}(p_{2},p_{1})\right),$
$=$ $\log Z^{G}(p_{1},p_{2})+\frac{1}{4}D_{J}(p_{1},p_{2}),$
$=$ $\frac{1}{4}D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}).$
$\square$
Corollary 1 (G-JSD upper bound)
We have the upper bound $D_{\mathrm{JS}}^{G}(p,q)\leq\frac{1}{4}D_{J}(p,q)$.
Proof:
Since $D_{B}\geq 0$ and $D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2})$, we have $D_{\mathrm{JS}}^{G}(p,q)\leq\frac{1}{4}D_{J}(p,q)$. $\square$
We have $\lim_{\gamma\rightarrow-\infty}P_{\gamma}(a,b)=\min(a,b)$ and $\lim_{\gamma\rightarrow+\infty}P_{\gamma}(a,b)=\max(a,b)$. Thus we can define both (extented) $\min$-JSD and (extented) $\max$-JSD. Using the fact that $\min(a,b)=\frac{a+b}{2}-\frac{1}{2}|a-b|$ and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|a-b|$, we obtain the extremal mixture normalization terms as follows:
$Z_{\min}(p_{1},p_{2})$ $=$ $\int\min(p_{1},p_{2})\,{\mathrm{d}\mu}=1-D_{\mathrm{TV}}(p_{1},p_{2}),$ (12)
$Z_{\max}(p_{1},p_{2})$ $=$ $\int\max(p_{1},p_{2})\,{\mathrm{d}\mu}=1+D_{\mathrm{TV}}(p_{1},p_{2}),$ (13)
where $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$ is the total variation distance.
Proposition 4 ($\max$-JSD)
The following upper bound holds for $\max$-JSD:
$0\leq D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2}),$ (14)
where $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\,\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$ denotes the total variation distance. Furthermore, the following identity relates the two types of $\max$-JSDs:
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})=D_{\mathrm{JS}}^{\widetilde{\max}}(p_{1},p_{2})+D_{\mathrm{TV}}(p_{1},p_{2})-\log D_{\mathrm{TV}}(p_{1},p_{2}).$ (15)
Proof:
We have
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2}){:=}\frac{1}{2}\,\int\left(p_{1}\log\frac{p_{1}}{\max(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\max(p_{1},p_{2})}+2\max(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu}.$
Since both $\log\frac{p}{\max(p,q)}\leq 0$ and $\log\frac{q}{\max(p,q)}\leq 0$, and $\max(a,b)=\frac{a+b}{2}+\frac{1}{2}|b-a|$, we have
$D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq\int\left(\frac{p+q}{2}+\frac{1}{2}|q-p|-\frac{p+q}{2}\right)\,{\mathrm{d}\mu}.$
That is, $D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2})$ where $D_{\mathrm{TV}}(p_{1},p_{2})$ is the total variation distance.
We characterize the gap as follows:
$\Delta_{\max}(p_{1},p_{2})$ $=$ $Z_{\max}(p_{1},p_{2})-\log Z_{\max}(p_{1},p_{2})-1,$
$=$ $D_{\mathrm{TV}}(p_{1},p_{2})-\log(1+D_{\mathrm{TV}}(p_{1},p_{2}))\geq 0,$
since $0\leq D_{\mathrm{TV}}\leq 1$. Thus $D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{\max}}(p_{1},p_{2})$. $\square$
Proposition 5 (min-JSD)
We have the following lower bound on the extended min-JSD:
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}D_{J}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2}),$
where $D_{J}(p_{1},p_{2}){:=}D_{\mathrm{KL}}(p_{1},p_{2})+D_{\mathrm{KL}}(p_{2},p_{1})=\int(p_{1}-p_{2})\log\frac{p_{1}}{p_{2}}\,{\mathrm{d}\mu}$ is Jeffreys’ divergence [14, 28] and
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})=D_{\mathrm{JS}}^{{\min}}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2})+\log(1-D_{\mathrm{TV}}(p_{1},p_{2})).$
Proof:
We have
$\Delta_{\min}(p_{1},p_{2})$ $=$ $Z_{\min}(p_{1},p_{2})-\log Z_{\min}(p_{1},p_{2})-1,$
$=$ $-D_{\mathrm{TV}}(p_{1},p_{2})-\log(1-D_{\mathrm{TV}}(p_{1},p_{2}))\geq 0,$
since $0\leq D_{\mathrm{TV}}\leq 1$. Note that the gap can be arbitrarily large when $D_{\mathrm{TV}}(p_{1},p_{2})\rightarrow 1^{-}$.
Thus we have $D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq D_{\mathrm{JS}}^{{\min}}(p_{1},p_{2})$.
To get the lower bound, we use the fact that $\min(p_{1},p_{2})\leq\sqrt{p_{1}p_{2}}$. Indeed, we have
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})$ $=$ $\frac{1}{2}\left(\int(p_{1}\log\frac{p_{1}}{\min(p_{1},p_{2})}+p_{2}\log\frac{p_{2}}{\min(p_{1},p_{2})}+2\min(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu},$
$\geq$ $\frac{1}{2}\int\left(\frac{1}{2}p_{1}\log\frac{p_{1}}{p_{2}}+\frac{1}{2}p_{2}\log\frac{p_{2}}{p_{1}}+2\min(p_{1},p_{2})-(p_{1}+p_{2})\right)\,{\mathrm{d}\mu},$
$\geq$ $\frac{1}{4}D_{J}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2}).$
$\square$
Remark 4
Let us report the total variation distance between two univariate Gaussian distributions $p_{\mu_{1},\sigma_{1}}$ and $p_{\mu_{2},\sigma_{2}}$ in closed-form using the error function [24]: $\mathrm{erf}(x)=\frac{1}{\sqrt{\pi}}\int_{-x}^{x}e^{-t^{2}}\mathrm{d}t$.
  • When $\sigma_{1}=\sigma_{2}=\sigma_{2}$, we have
    $\mathrm{TV}(p_{1},p_{2})=\frac{1}{2}\left|\Phi(x^{*};\mu_{2},\sigma_{2})-\Phi(x^{*};\mu_{1},\sigma_{1})\right|,$ (16)
    where $\Phi(x;\mu,\sigma)=\frac{1}{2}(1+\mathrm{erf}(\frac{x-\mu}{\sigma\sqrt{2}}))$ is the cumulative distribution, and
    $x^{*}=\frac{\mu_{1}^{2}-\mu_{2}^{2}-2\log\frac{\sigma_{2}}{\sigma}}{2(\mu_{1}-\mu_{2})}.$ (17)
  • When $\sigma_{1}\not=\sigma_{2}$, we let $x_{1}=\frac{-b-\sqrt{\Delta}}{2a}$ and $x_{2}=\frac{-b+\sqrt{\Delta}}{2a}$ where $\Delta=b^{2}-4ac\geq 0$ and
    $a$ $=$ $\frac{1}{\sigma_{1}^{2}}-\frac{1}{\sigma_{2}^{2}},$ (18)
    $b$ $=$ $2\left(\frac{\mu_{2}}{\sigma_{2}}-\frac{\mu_{1}}{\sigma_{1}}\right),$ (19)
    $c$ $=$ $\left(\frac{\mu_{1}}{\sigma_{1}}\right)^{2}-\left(\frac{\mu_{2}}{\sigma_{2}}\right)^{2}-2\log\frac{\sigma_{2}}{\sigma_{1}}.$ (20)
    The total variation is given by
    $\mathrm{TV}(p_{1},p_{2})=$ (21)
    $\frac{1}{2}\left(\left|\mathrm{erf}\left(\frac{x_{1}-\mu_{1}}{\sigma_{1}\sqrt{2}}\right)-\mathrm{erf}\left(\frac{x_{1}-\mu_{2}}{\sigma_{2}\sqrt{2}}\right)\right|+\left|\mathrm{erf}\left(\frac{x_{2}-\mu_{1}}{\sigma_{1}\sqrt{2}}\right)-\mathrm{erf}\left(\frac{x_{2}-\mu_{2}}{\sigma_{2}\sqrt{2}}\right)\right|\right)$
Next, we shall consider the important case of $p_{1}$ and $p_{2}$ belonging to the family of multivariate normal distributions, commonly called Gaussian distributions.
3 Geometric JSDs between Gaussian distributions
3.1 Exponential families
The formula for the G-JSD between two Gaussian distributions was reported in [25] using the more general framework of exponential families. An exponential family [3] is a family of probability measures $\{P_{\lambda}\}$ with Radon-Nikodym densities $p_{\lambda}$ with respect to $\mu$ expressed canonically as
$p_{\lambda}(x)$ ${:=}$ $\exp\left(\left\langle\theta(\lambda),t(x)\right\rangle-F(\theta)+k(x)\right),$
$=$ $\frac{1}{Z(\theta)}\,\exp\left(\left\langle\theta(\lambda),t(x)\right\rangle+k(x)\right),$
where $\theta(\lambda)$ is the natural parameter, $t(x)$ the sufficient statistic, $k(x)$ an auxiliary carrier term with respect to $\mu$, and $F(\theta)$ the cumulant function. The partition function $Z(\theta)$ is the normalizer denominator: $Z(\theta)=\exp(F(\theta))$. The cumulant function $F(\theta)=\log Z(\theta)$ is strictly convex and analytic [3], and the partition function $Z(\theta)=\exp(F(\theta))$ is strictly log-convex (and hence also strictly convex).
We consider the exponential family of multivariate Gaussian distributions
$\mathcal{N}=\{N(\mu,\Sigma){\ :\ }\mu\in\mathbb{R}^{d},\Sigma\in\mathrm{PD}(d)\},$
where $\mathrm{PD}(d)$ denotes the set of symmetric positive-definite matrices of size $d\times d$. Let $\lambda{:=}(\lambda_{v},\lambda_{M})=(\mu,\Sigma)$ denote the compound (vector,matrix) parameter of a Gaussian. The $d$-variate Gaussian density is given by
$p_{\lambda}(x;\lambda)$ ${:=}$ $\frac{1}{(2\pi)^{\frac{d}{2}}\sqrt{|\lambda_{M}|}}\exp\left(-\frac{1}{2}(x-\lambda_{v})^{\top}\lambda_{M}^{-1}(x-\lambda_{v})\right),$ (22)
where $|\cdot|$ denotes the matrix determinant. The natural parameters $\theta$ are expressed using both a vector parameter $\theta_{v}$ and a matrix parameter $\theta_{M}$ in a compound parameter $\theta=(\theta_{v},\theta_{M})$. By defining the following compound inner product on a compound (vector,matrix) parameter
$\left\langle\theta,\theta^{\prime}\right\rangle{:=}\theta_{v}^{\top}\theta_{v}^{\prime}+\mathrm{tr}\left({\theta_{M}^{\prime}}^{\top}\theta_{M}\right),$ (23)
where $\mathrm{tr}(\cdot)$ denotes the matrix trace, we rewrite the Gaussian density of Eq. 22 in the canonical form of an exponential family:
$p_{\theta}(x;\theta)$ ${:=}$ $\exp\left(\left\langle t(x),\theta\right\rangle-F_{\theta}(\theta)\right)=p_{\lambda}(x),$ (24)
where $\theta=\theta(\lambda)$ with
$\theta=(\theta_{v},\theta_{M})=\left(\Sigma^{-1}\mu,-\frac{1}{2}\Sigma^{-1}\right)=\theta(\lambda)=\left(\lambda_{M}^{-1}\lambda_{v},-\frac{1}{2}\lambda_{M}^{-1}\right),$ (25)
is the compound vector-matrix natural parameter and
$t(x)=(x,-xx^{\top})$ (26)
is the compound vector-matrix sufficient statistic. There is no auxiliary carrier term (i.e., $k(x)=0$). The function $F_{\theta}$ is given by:
$F_{\theta}(\theta)=\frac{1}{2}\left(d\log\pi-\log|\theta_{M}|+\frac{1}{2}\theta_{v}^{\top}\theta_{M}^{-1}\theta_{v}\right),$ (27)
Remark 5
Beware that when the cumulant function is expressed using the ordinary parameter $\lambda=(\mu,\Sigma)$, the cumulant function $F_{\theta}(\theta(\lambda))$ is not anymore convex:
$F_{\lambda}(\lambda)$ $=$ $\frac{1}{2}\left(\lambda_{v}^{\top}\lambda_{M}^{-1}\lambda_{v}+\log|\lambda_{M}|+d\log 2\pi\right),$ (28)
$=$ $\frac{1}{2}\left(\mu^{\top}\Sigma^{-1}\mu+\log|\Sigma|+d\log 2\pi\right).$ (29)
We convert between the ordinary parameterization $\lambda=(\mu,\Sigma)$ and the natural parameterization $\theta$ using these formula:
$\theta=(\theta_{v},\theta_{M})=\left\{\begin{array}[]{ll}\theta_{v}(\lambda)=\lambda_{M}^{-1}\lambda_{v}=\Sigma^{-1}\mu\\ \theta_{M}(\lambda)=\frac{1}{2}\lambda_{M}^{-1}=\frac{1}{2}\Sigma^{-1}\end{array}\right$. $\Leftrightarrow$ $\lambda=(\lambda_{v},\lambda_{M})=\left\{\begin{array}[]{ll}\lambda_{v}(\theta)=\frac{1}{2}\theta_{M}^{-1}\theta_{v}=\mu\\ \lambda_{M}(\theta)=\frac{1}{2}\theta_{M}^{-1}=\Sigma\end{array}\right$. (34)
The geometric mixture $p_{\theta_{1}}^{\alpha}p_{\theta_{2}}^{1-\alpha}$ of two densities of an exponential family is a density $p_{\alpha\theta_{1}+(1-\alpha)\theta_{2}}$ of the exponential family with partition function $Z_{\alpha}(\theta_{1},\theta_{2})=\exp(-J_{F,\alpha}(\theta_{1},\theta_{2}))$ where $J_{F,\alpha}(\theta_{1},\theta_{2})$ denotes the skew Jensen divergence [17, 27]:
$J_{F,\alpha}(\theta_{1},\theta_{2}){:=}\alpha F(\theta_{1})+(1-\alpha)F(\theta_{2})-F(\alpha\theta_{1}+(1-\alpha)\theta_{2}).$
Therefore the difference gap of Eq. 10 between the G-JSD and the extended G-JSD between exponential family densities is given by:
$\Delta(\theta_{1},\theta_{2})$ $=$ $\exp({-J_{F,\alpha}(\theta_{1},\theta_{2})})+J_{F,\alpha}(\theta_{1},\theta_{2})-1,$ (36)
$=$ $Z_{\alpha}(\theta_{1},\theta_{2})-\log Z_{\alpha}(\theta_{1},\theta_{2})-1,$ (37)
$=$ $Z_{\alpha}(\theta_{1},\theta_{2})-F(\alpha\theta_{1}+(1-\alpha)\theta_{2})-1.$ (38)
Since $Z_{\alpha}=\exp(-J_{F,\alpha}(\theta_{1},\theta_{2}))\leq 1$, the gap $\Delta$ is negative, and we have
$D_{\mathrm{JS}^{+}}^{\tilde{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}})\leq D_{\mathrm{JS}}^{{G}_{\alpha},\beta}(p_{\mu_{1},\Sigma_{1}},p_{\mu_{2},\Sigma_{2}}).$
Corollary 2
When $p_{1}=p_{\theta_{1}}$ and $p_{2}=p_{\theta_{2}}$ belongs to a same exponential family with cumulant function $F(\theta)$, we have
$D_{\mathrm{JS}}^{G}(p_{\theta_{1}},p_{\theta_{2}})={\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (39)
since $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle$ amounts to a symmetrized Bregman divergence [28].
Proof:
We have $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))$ and $D_{J}(p_{\theta_{1}},p_{\theta_{2}})=J_{F}(\theta_{1},\theta_{2})$. $\square$
The extended geometric Jensen–Shannon divergence and geometric Jensen–Shannon divergence between two densities of an exponential family is given by
$D_{\mathrm{JS}}^{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ ${\frac{1}{4}(\theta_{2}-\theta_{1})^{\top}(\nabla F(\theta_{2})-\nabla F(\theta_{1}))}-{\left(\frac{F(\theta_{1})+F(\theta_{2})}{2}-F\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\right)},$ (40)
$D_{\mathrm{JS}}^{\tilde{G}}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $\frac{1}{4}\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})-\nabla F(\theta_{1})\right\rangle-\exp(-J_{F}(\theta_{1},\theta_{2}))-1,$ (41)
$D_{\mathrm{JS}^{*}}^{G}(p_{\theta_{1}},p_{\theta_{2}})$ $=$ $J_{F}(\theta_{1},\theta_{2})$ (42)
3.2 Closed-form formula for Gaussian distributions
Let us report the corresponding closed-form formula for Gaussian distributions:
Corollary 3
The skew $G$-Jensen–Shannon divergence $\mathrm{JS}_{\alpha}^{G}$ and the dual skew $G$-Jensen–Shannon divergence ${\mathrm{JS}^{*}}_{\alpha}^{G}$ between two $d$-variate Gaussian distributions $N(\mu_{1},\Sigma_{1})$ and $N(\mu_{2},\Sigma_{2})$ is
${\mathrm{JS}}^{G_{\alpha}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $(1-\alpha)\,D_{\mathrm{KL}}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{\alpha},\Sigma_{\alpha})})+\alpha\,D_{\mathrm{KL}}(p_{(\mu_{2},\Sigma_{2})},p_{(\mu_{\alpha},\Sigma_{\alpha})}),$
$=$ $(1-\alpha)\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{1})+\alpha\,B_{F}((\theta_{1}\theta_{2})_{\alpha},\theta_{2}),$
$=$ $\frac{1}{2}\left(\mathrm{tr}\left(\Sigma_{\alpha}^{-1}((1-\alpha)\Sigma_{1}+\alpha\Sigma_{2})\right)+\log\left(\frac{|\Sigma_{\alpha}|}{|\Sigma_{1}|^{1-\alpha}|\Sigma_{2}|^{\alpha}}\right)\right$.
$=$ $\left.+(1-\alpha)(\mu_{\alpha}-\mu_{1})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{1})+\alpha(\mu_{\alpha}-\mu_{2})^{\top}\Sigma_{\alpha}^{-1}(\mu_{\alpha}-\mu_{2})-d\right)$
${\mathrm{JS}}^{G_{\alpha}}_{*}(p_{(\mu_{1},\Sigma_{1})},p_{(\mu_{2},\Sigma_{2})})$ $=$ $(1-\alpha)\,D_{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{1},\Sigma_{1})})+\alpha D_{\mathrm{KL}}(p_{(\mu_{\alpha},\Sigma_{\alpha})},p_{(\mu_{2},\Sigma_{2})}),$
$=$ $(1-\alpha)\,B_{F}(\theta_{1},(\theta_{1}\theta_{2})_{\alpha})+\alpha\,B_{F}(\theta_{2},(\theta_{1}\theta_{2})_{\alpha}),$
$=$ $J_{F,\alpha}(\theta_{1},\theta_{2}),$
$=$ $\frac{1}{2}\left((1-\alpha)\mu_{1}^{\top}\Sigma_{1}^{-1}\mu_{1}+\alpha\mu_{2}^{\top}\Sigma_{2}^{-1}\mu_{2}-\mu_{\alpha}^{\top}\Sigma_{\alpha}^{-1}\mu_{\alpha}+\log\frac{|\Sigma_{1}|^{1-\alpha}|\Sigma_{2}|^{\alpha}}{|\Sigma_{\alpha}|}\right),,$
$F(\mu,\Sigma)$ $=$ $\frac{1}{2}\left(\mu^{\top}\Sigma^{-1}\mu+\log|\Sigma|+d\log 2\pi\right),$
$F(\theta_{v},\theta_{M})$ $=$ $\frac{1}{2}\left(d\log\pi-\log|\theta_{M}|+\frac{1}{2}\theta_{v}^{\top}\theta_{M}^{-1}\theta_{v}\right),$
$\Delta(\theta_{1},\theta_{2})$ $=$ $\exp({-J_{F,\alpha}(\theta_{1},\theta_{2})})+J_{F,\alpha}(\theta_{1},\theta_{2})-1,$
where $\Sigma_{\alpha}$ is the matrix harmonic barycenter:
$\Sigma_{\alpha}=\left((1-\alpha)\Sigma_{1}^{-1}+\alpha\Sigma_{2}^{-1}\right)^{-1},$ (43)
and
$\mu_{\alpha}=\Sigma_{\alpha}\left((1-\alpha)\Sigma_{1}^{-1}\mu_{1}+\alpha\Sigma_{2}^{-1}\mu_{2}\right).$ (44)
4 Extended and normalized G-JSDs as regularizations of the JSD
The $M$-Jensen–Shannon divergence $\mathrm{JS}_{\mathrm{KL}}^{M}(p,q)$ can be interpreted as a regularization of the ordinary JSD:
Proposition 6 (JSD regularization)
For any arbitrary mean $M$, the following identity holds:
$D_{\mathrm{JS}}^{M}(p_{1},p_{2})=D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right).$ (45)
Notice that $(p_{1}p_{2})_{A}=\frac{p_{1}+p_{2}}{2}$.
Proof:
We have
$\mathrm{JS}_{\mathrm{KL}}^{M}(p_{1},p_{2})$ ${:=}$ $\frac{1}{2}\left(D_{\mathrm{KL}}(p_{1}:(p_{1}p_{2})_{M})+D_{\mathrm{KL}}(p_{2}:(p_{1}p_{2})_{M})\right),$
$=$ $\frac{1}{2}\int\left(p_{1}\log\frac{p_{1}\,(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}\,(p_{1}p_{2})^{A}}+p_{2}\log\frac{p_{2}\,(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}\,(p_{1}p_{2})_{A}}\right)\,{\mathrm{d}\mu},$
$=$ $\frac{1}{2}\int\left(p_{1}\log\frac{p_{1}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}+p_{2}\log\frac{p_{2}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\right)\,{\mathrm{d}\mu},$
$=$ $\frac{1}{2}\int\left(p_{1}\log\frac{p_{1}}{(p_{1}p_{2})_{A}}+p_{2}\log\frac{p_{2}}{(p_{1}p_{2})_{A}}\right)\,{\mathrm{d}\mu}+\int\frac{1}{2}\,(p_{1}+p_{2})\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\,{\mathrm{d}\mu},$
$=$ $D_{\mathrm{JS}}(p_{1},p_{2})+\int(p_{1}p_{2})_{A}\log\frac{(p_{1}p_{2})_{A}}{(p_{1}p_{2})_{M}}\,{\mathrm{d}\mu},$
$=$ $D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}).$
$\square$
Corollary 4 ($M$-JSD lower bound)
We have $D_{\mathrm{JS}}^{M}(p,q)\geq D_{\mathrm{JS}}(p,q)$.
Proof:
Since $D_{\mathrm{JS}}^{M}(p,q)=D_{\mathrm{JS}}(p,q)+D_{\mathrm{KL}}\left(\frac{p+q}{2},(pq)_{M}\right)$ and $D_{\mathrm{KL}}\geq 0$ by Gibbs’ inequality, we have $D_{\mathrm{JS}}^{M}(p,q)\geq D_{\mathrm{JS}}(p,q)$. $\square$
Since the extended $M$-JSD is $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}^{{M}_{\alpha},\beta}(p_{1},p_{2})+Z_{\alpha}-\log(Z_{\alpha})-1$, the extended $M$-JSD $D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}$ can also be interpreted as another regularization of the Jensen–Shannon divergence when dealt with probability densities:
$D_{\mathrm{JS}^{+}}^{\tilde{M}_{\alpha},\beta}(p_{1},p_{2})=D_{\mathrm{JS}}(p_{1},p_{2})+D_{\mathrm{KL}}\left(\frac{p_{1}+p_{2}}{2},(p_{1}p_{2})_{M}\right)+Z_{\alpha}(p_{1},p_{2})-\log(Z_{\alpha}(p_{1},p_{2}))-1.$ (46)
5 Summary and concluding remarks
In this paper, we first recalled the Jensen–Shannon symmetrization (JS-symmetrization) scheme of [25] for an arbitrary statistical dissimilarity $D(\cdot,\cdot)$ using an arbitrary weighted scalar mean $M_{\alpha}$ as follows:
$\mathrm{JS}_{D}^{M_{\alpha},\beta}(p_{1},p_{2}){:=}\beta\,D\left(p_{1},{(p_{1}p_{2})}_{\alpha}^{M}\right)+(1-\beta)\,D\left(p_{2},{(p_{1}p_{2})}_{\alpha}^{M}\right),\quad(\alpha,\beta)\in(0,1)^{2},$
In particular, we showed that the skewed Bhattacharyya distance and the Chernoff information can both be interpreted as JS-symmetrizations of the reverse Kullback–Leibler divergence.
Then we defined two types of geometric Jensen–Shannon divergence between probability densities: The first type $D_{\mathrm{JS}}^{M}$ requires to normalize $M$-mixtures and relies on the Kullback–Leibler divergence: $D_{\mathrm{JS}}^{M}=\mathrm{JS}_{\mathrm{KL}}^{M_{\frac{1}{2}},\frac{1}{2}}$. The second type $D_{\mathrm{JS}^{+}}^{\tilde{M}}$ does not normalize $M$-mixtures and uses the extended Kullback–Leibler divergence: $D_{\mathrm{JS}^{+}}^{\tilde{M}}=\mathrm{JS}_{{\mathrm{KL}}^{+}}^{M_{\frac{1}{2}},\frac{1}{2}}$. When $M$ is the arithmetic mean $A$, both $M$-JSD types coincide with the ordinary Jensen–Shannon divergence of Eq. 2.
We have shown that that both geometric JSD types can be interpreted as regularized Jensen–Shannon divergences $D_{\mathrm{JS}}$ by additive terms:
$D_{\mathrm{JS}}^{M}(p_{1},p_{2})$ $=$ ${\bf D_{\mathrm{JS}}(p_{1},p_{2})}+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M}),$
$D_{\mathrm{JS}^{+}}^{\tilde{M}}(p_{1},p_{2})$ $=$ $D_{\mathrm{JS}}^{M}(p_{1},p_{2})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
$=$ ${\bf D_{\mathrm{JS}}(p_{1},p_{2})}+D_{\mathrm{KL}}((p_{1}p_{2})_{A},(p_{1}p_{2})_{M})+Z_{M}(p_{1},p_{2})-\log Z_{M}(p_{1},p_{2})-1,$
where $Z_{M}(p_{1},p_{2})=\int M(p_{1},p_{2})\,{\mathrm{d}\mu}$ is the $M$-mixture normalizer.
We have expressed $D_{\mathrm{JS}}^{G}$ using both Jeffreys $D_{J}$ and Bhattacharrya $D_{B}$ divergences as follows:
$D_{\mathrm{JS}}^{G}(p_{1},p_{2})=\frac{1}{4}\,D_{J}(p_{1},p_{2})-D_{B}(p_{1},p_{2}).$
We obtained both lower and upper bounds on $D_{\mathrm{JS}}^{G}$:
$0\leq D_{\mathrm{JS}}(p_{1},p_{2})\leq D_{\mathrm{JS}}^{G}(p_{1},p_{2})\leq\frac{1}{4}\,D_{J}(p_{1},p_{2}).$
We studied power JSDs by considering the power means and study in the $\pm\infty$ limits, the extended max-JSD and min-JSD: We proved that the extended max-JSD is upper bounded by the total variation distance $D_{\mathrm{TV}}(p_{1},p_{2})=\frac{1}{2}\,\int|p_{1}-p_{2}|\,{\mathrm{d}\mu}$:
$0\leq D_{\mathrm{JS}^{+}}^{\widetilde{\max}}(p_{1},p_{2})\leq D_{\mathrm{TV}}(p_{1},p_{2}),$
and that the extended min-JSD is lower bounded as follows:
$D_{\mathrm{JS}^{+}}^{\widetilde{\min}}(p_{1},p_{2})\geq\frac{1}{4}D_{J}(p_{1},p_{2})-D_{\mathrm{TV}}(p_{1},p_{2}),$
where $D_{J}$ denotes Jeffreys’s divergence: $D_{J}(p_{1},p_{2})=D_{\mathrm{KL}}(p_{1},p_{2})+D_{\mathrm{KL}}(p_{2},p_{1})$.
The advantage of using the extended G-JSD is that we do not need to normalize geometric mixtures while this novel divergence retains the property that it amounts to a regularization of the ordinary Jensen–Shannon divergence by an extra additive gap term.
Finally, we expressed $D_{\mathrm{JS}}^{G}$ (Eq. 40) and $D_{\mathrm{JS}^{+}}^{\tilde{G}}$ (Eq. 41) for exponential families, characterized the gap between these two types of divergences as a function of the cumulant and partition functions, and reported corresponding explicit formula for the multivariate Gaussian (exponential) family. The G-JSD between Gaussian distributions has already been used successfully in many applications [8, 20, 18, 23, 33, 39, 34, 37, 13].
As a final word, let us also mention a recent work that considered the geometric Jensen–Shannon divergence between Gaussian measures in infinite-dimensional Hilbert spaces [32] (with applications to Gaussian processes left for future work).